# TopoBenchmarkX: A Framework for Benchmarking Topological Deep Learning

### v1

@@ 4.2 Topological liftings
This section explores different (structural) liftings implemented in TopoBenchmarkX. In particular, liftings of graphs to cell complexes, simplicial complexes, and hypergraphs are discussed. [^3] Note that the modular design of TopoBenchmarkX easily allows for the addition of new lifting transforms between any pair of topological domains.
From graphs to cell complexes: cycle-based liftings. A graph is lifted to a cell complex as follows. First, a finite set of cycles (closed loops) within the graph is identified. Second, each identified cycle in the graph is associated with a 2-cell, such that the boundary of the 2-cell is the cycle. The nodes and edges of the cell complex are inherited from the graph.
From graphs to simplicial complexes: clique complexes. By lifting a graph to a simplicial complex, both pairwise and higher-order interactions are captured. For a graph, the corresponding clique complex is formed by considering every complete subgraph (clique) as a simplex. Specifically, each node is a 0-simplex, each edge (clique of size 2) is a 1-simplex, each triangle (clique of size 3) is a 2-simplex, and so forth. In general, a clique of size $k+1$ becomes a $k$-simplex.
From graphs to simplicial complexes: neighbor complexes. Neighbor complexes lift the neighborhoods of nodes to become simplices as follows. For each node in the graph, take the set of its neighbors and the node itself. These sets are then treated as simplices. The dimension of each simplex depends on the degree of the nodes. For example, a node with $d$ neighbors forms a $d$-simplex.
From graphs to hypergraphs: $k$-hop liftings. Let $\mathcal{G}=(V,E)$ be a graph and $\mathcal{H}=(V,\mathcal{E})$ be a hypergraph. The $k$-neighborhood $N_{k}(v)$ of a node $v\in V$ in $\mathcal{G}$ consists of all nodes reachable within $k$ steps from $v$. To lift $\mathcal{G}$ to $\mathcal{H}$, for each node $v\in V$, determine $N_{k}(v)$ and create a hyperedge $e_{v}$ in $\mathcal{H}$ such that $e_{v}=N_{k}(v)$. Thus, the set of hyperedges in $\mathcal{H}$ is given by =$\mathcal{E}=\{N_{k}(v)\mid v\in V\}$.
@@ 5 Numerical experiments
This section presents numerical experiments that exemplify the wide range of functionality of TopoBenchmarkX, and provide a first attempt to perform a broad cross-domain comparison. First, the general experimental setup is described. The experiments compare twelve graph, hypergraph and topological neural network models across four learning tasks related to twenty-two datasets. Subsequently, the main benchmark results are presented. Finally, it is shown how TopoBenchmarkX facilitates comparisons in TDL by conducting an ablation study for different types of signal propagation [^4] See Appendix E for an introduction to TNNs and higher-order message passing on topological domains..
@@ 5.1 Setup
Learning tasks and datasets. Four learning tasks are carried out, namely node classification (seven datasets), node regression (seven datasets), graph classification (seven datasets), and graph regression (one dataset). For node classification, cocitation datasets (Cora, Citeseer, and Pubmed), as well as heterophilic datasets (Amazon rating, Minesweeper, Roman empire, and Tolokers) are used (Platonov et al., 2023). For node regression, the election, bachelor, birth, death, income, migration, and unemployment datasets from US election map networks are adapted (Jia and Benson, 2020). In these datasets, nodes represent US states, edges connect neighboring states, and each state is characterized by demographic and election statistics. For each dataset, one statistic is set to be the output variable, while the remaining statistics serve as node features, referring to each such dataset with the output statistic’s name. For graph classication, the TUDataset is used, consisting of the MUTAG, PROTEINS, NCI1, NCI109, IMDB-BIN, IMDB-MUL, and REDDIT datasets; see Morris et al. (2020). For graph regression, the ZINC dataset is employed (Irwin et al., 2012). See Table 7 in Appendix F for descriptive summaries of the datasets.
@@
Models. The twelve neural network models are supported on four domains: graphs, hypergraphs, simplicial complexes, and cell complexes. In particular, three GNNs (GCN, GIN, and GAT), three hypergraph neural networks (EDGNN, AllSetTransformer, and UniGNN2), three simplicial neural networks (SCN, SCCN, and SCCNN), and three cell complex neural networks (CCXN, CWN, and CCCN) are used in the experiments. Please see Appendix E.2 for further details about these architectures, including a comparison of their hyperparameter budget.
@@
Training and evaluation. Five splits are generated for each dataset. Each split comprises 50%/25%/25% of data points in the training, validation, and test set, respectively. As an exception, the pre-defined splits for the ZINC dataset are used (Irwin et al., 2012). The optimal hyperparameter set is selected based on the best average performance over five validation splits. See Appendix C.2 for more details on the hyperparameter search. One performance metric per dataset is reported. More specifically, the predictive accuuracy is reported for Cora, Citeseer, Pubmed, Amazon, Empire, MUTAG, PROTEINS, NCI1, NCI109, IMDB-BIN, IMDB-MUL, and REDDIT; the AUC-ROC metric is reported for Minesweeper and Tolokers; the mean-squared error is reported for election, bachelor, birth, death, income, migration, and unemployment; and the mean-absolute error is reported for ZINC. For each performance metric, its mean and standard deviation are computed over the five test sets per dataset and are reported in Table 1 (note that OOM stands for ‘out of memory’ in Table 1).
@@ 5.2 Main results
As seen from Table 1, higher-order neural networks (based on hypergraphs, simplicial and cell complexes) achieve the best performance on sixteen of twenty-two datasets, whereas GNNs achieve the best performance on seven datasets. GNNs perform best on node regression in the majority of cases (five out of seven). These best results obtained by GNNs are closely matched by TNNs, since the latter attain metrics within one standard deviation from the former. In contrast, in nine out of sixteen datasets TNNs outperform GNNs, and attain performance metrics that are higher by more than one standard deviation with respect to GNNs. In other words, in the cases where higher-order networks outperform GNNs, the performance gap is more pronounced. It is also noted that, for demonstration purposes, only one fixed lifting is considered to transform graph data to each of the considered topological domains (see Appendix C.3). These results suggest that TNNs hold an advantage over GNNs, even without lifting optimization. Most importantly for the context of this paper, the benchmarks demonstrate the extent of comparisons that can be performed with TopoBenchmarkX across models and datasets.
@@ 5.3 Ablation study
Graph and hypergraph (neural network) models differ from simplicial and cell complex (neural network) models in terms of the domains and representations they support. Graph and hypergraph models can output two types of representations: node representations and edge or hyperedge representations. In contrast, the output of simplicial and cell models depends on the different types of cells present (00-cell up to $n$-cells) and the model itself. For example, a simplicial or cell complex model may process an $n$-cell input but may not produce an $n$-cell output. The backbone_wrapper in TopoBenchmarkX addresses these differences in the underlying domains of the models.
@@
There is a second difference, which is inherent in the TNNs themselves. Consider a downstream classification task. For graphs, the standard practice is to perform classification over pooled node features. However, this aspect has not been extensively studied in the TDL literature. For instance, a simplicial or cell model may update 1-cell representations (edges) or 2-cell representations (cycles or triangles) while leaving 0-cell (nodes) representations unchanged, making direct pooling over nodes potentially meaningless. One can consider more complicated update processes in which different $n$-cell representations are combined, which makes the notion of pooling more intricate for higher-order domains, in general. These architectural implications are complex and are related to open research questions in TDL.
@@
However, to compare different neural network architectures, the second aforementioned difference must be addressed. To this end, the ablation study of this section considers two types of readouts to perform a rigorous evaluation: direct readout (DR), in which the downstream task is performed directly over the 00-cell representation, and signal down-propagation (SDP), in which information from higher-order cell representations is iteratively fused down to 00-cell representations using appropriate incidence matrices, followed by a linear projection over the concatenated $n-1$-cell signal and the fused $n-1$-cell representation. For instance, if a simplicial or cell complex model outputs 00-cell, $1$-cell, and $2$-cell representations, the signal propagates from $2$-cells to $1$-cells, and from $1$-cells to 00-cells during readout. The downstream task is then performed over the 00-cell representations. [^5] See Appendix E for an introduction to TNNs and higher-order message passing on topological domains.
@@
Table 2 shows that the best-performing readout type depends on the model’s interior signal propagation. The CWN model does not update the 00-cell representation during propagation and the SDP readout strategy performs notably better, as expected. Conversely, CCCN, SCCNN, and SCN propagate information to 00-cells, so incorporating SDP readout leads to either small or minor variations in performance metrics. Please refer to Appendix C.3 for more detailed results.
@@
These results highlight the importance of considering the structural properties of TNNs. The performance variations observed in this ablation study emphasize how critical architectural and lifting decisions are in higher-order learning models. By facilitating comparisons across a wide range of models and datasets, TopoBenchmarkX enables researchers to derive new insights and advance TDL.
@@ 6 Concluding remarks, limitations, and future work
This paper has introduced TopoBenchmarkX, an open-source benchmarking framework for TDL. TopoBenchmarkX simplifies the benchmarking process and accelerates research by organizing the TDL pipeline into a sequence of modular steps. A key feature of TopoBenchmarkX is its ability to lift between topological domains, allowing richer data representations and more detailed analyses. This capability supports the mapping of graph topology and features to higher-order topological domains, such as simplicial and cell complexes. The effectiveness and applicability of TopoBenchmarkX has been illustrated by benchmarking several TDL architectures on many learning tasks and datasets, gaining some preliminary insights about the relative advantages of different models.
The modular and scalable properties of TopoBenchmarkX open up several opportunities for future enhancements, and simultaneously allow us to address its current limitations. One area for future work is the integration of learnable liftings, not yet implemented in the framework. This is a promising direction, since learnable liftings can learn optimal representations of topological domains for specific downstream learning tasks. A second relevant limitation is the current lack of higher-order standard datasets. Thus, a second future work direction is to incorporate higher-order datasets into TopoBenchmarkX, giving users the option to choose between downloading built-in datasets (not yet available software feature) and generating synthetic higher-order data by lifting (already available software feature). Lastly, a third critical aspect pertains to the set of performance metrics. A wider range of alternative evaluation metrics specific to TDL needs to be developed and implemented to assess expressivity, explainability, and fairness (Papamarkou et al., 2024). More generally, users and researchers are encouraged to contribute to the development of TopoBenchmarkX in any of these three or other preferred directions.
Appendix A Societal impact and ethics
TopoBenchmarkX aims to standardize benchmarking and accelerate research in Topological Deep Learning, so we do not expect it to have any direct negative societal impact from its usage. Moreover, authors and contributors advocate for responsible use of TopoBenchmarkX, adherence to ethical guidelines, and promote transparency, fairness, and inclusivity in research.
The TopoBenchmarkX library will be constantly maintained to respect proprietary content. It will implement strict revision processes to ensure that all code implementations, libraries, and datasets have open-source licenses that guarantee their legitimate usage within the framework.
Appendix B Code availability and reproducibility
@@
The code for TopoBenchmarkX is available on GitHub under MIT license: https://github.com/pyt-team/TopoBenchmarkX. The codebase adheres to best software practices, employs continuous integration, is fully documented, and provides API documentation on its website https://pyt-team.github.io/topobenchmarkx/index.html.

### v2

@@ 4.3 Topological Liftings
In the context of TDL, as outlined in Section 2, liftings facilitate the mapping of data from one topological representation to another. This mapping comprises two key aspects: structural lifting and feature lifting (see Figure 3 for a visual example, and a formal definition can be found in Appendix A.2). Informally, the structural lifting is responsible for the transformation of the underlying relationships or elements of the data. For instance, it might determine how nodes and edges in a graph are mapped into triangles and tetrahedra in a simplicial complex. This structural transformation can be further categorized into connectivity-based, where the mapping relies solely on the existing connections within the data, and feature-based, where the data’s inherent properties or features guide or even fully determine the new structure. Feature lifting, conversely, addresses the transfer of data attributes or features during mapping, ensuring that the properties associated with the data elements are consistently preserved in the new representation, thus maintaining information integrity. Both structural and feature liftings are crucial for the effective application of TDL to diverse and complex datasets.
Figure 3: An illustration of lifting a graph (center) to two different topological domains: a simplicial complex (left) and a cell complex (right). The structural lifting maps the nodes and edges of the graph to higher-order topological structures, such as faces, while the feature lifting ensures the associated feature functions are consistently transferred between domains.
Table LABEL:tab:submissions in Appendix B provides a comprehensive list of all of the liftings currently implemented in TopoBench. Currently, TopoBench supports 11 structural liftings targeting simplicial complexes; 2 targeting cell complexes; 10 moving to hypergraphs; and 3 from a domain to combinatorial complexes. Notably, TopoBench’s modular design simplifies the integration of additional liftings, ensuring the framework’s adaptability to evolving research needs.
@@ 5 Numerical Experiments
This section presents numerical experiments that illustrate the breadth of TopoBench’s functionality by performing a cross-domain comparison. [^3] All code and instructions to reproduce the experiments of this section are publicly available; see Appendix Code Availability and Reproducibility for details on code availability and reproducibility. The overall setup is first described, then results from benchmarking various graph, hypergraph, and TNNs are reported, and an ablation study on signal propagation is presented to demonstrate how TopoBench supports comparisons in TDL.
@@ 5.1 Setup
Learning tasks and datasets. Four types of tasks are considered: node classification (seven datasets), node regression (seven datasets), graph classification (seven datasets), and graph regression (one dataset). For node classification, the cocitation datasets (Cora, Citeseer, and PubMed) and heterophilic datasets (Amazon Ratings, Minesweeper, Roman Empire, and Tolokers) are used (Platonov et al., 2023). For node regression, the election, bachelor, birth, death, income, migration, and unemployment datasets from US election map networks are adapted (Jia and Benson, 2020). In these datasets, each node represents a US state, edges connect neighboring states, and each state is characterized by demographic and election statistics. For each dataset, one statistic is designated as the target, while the others serve as node features, with the dataset named for the chosen target statistic. For graph classification, the TUDataset collection is used, specifically MUTAG, PROTEINS, NCI1, NCI109, IMDB-BIN, IMDB-MUL, and REDDIT (Morris et al., 2020). For graph regression, the ZINC dataset is employed (Irwin et al., 2012).
Higher-order datasets are constructed by lifting these graph datasets. For demonstration purposes, one structural lifting is considered for each of the considered higher-order topological spaces: cycle-based lifting for the cell domain (see Example 5), clique complex lifting for the simplicial domain (see Example 6), and $k$-hop lifting for the hypergraph domain (see Example 8). As for the feature lifting, the projected sum is always considered in all of these scenarios. Descriptive statistics for these topological versions of the datasets are provided in Table 4 in Appendix C.3.
@@
Models. Twelve neural networks, supported across four domains (graphs, hypergraphs, simplicial complexes, and cell complexes), are benchmarked. These include three GNNs (GCN, GIN, and GAT), three hypergraph neural networks (EDGNN, AllSetTransformer, and UniGNN2), three simplicial neural networks (SCN, SCCN, and SCCNN), and three cell complex neural networks (CCXN, CWN, and CCCN). Details on these architectures and their hyperparameters appear in Appendix C, while the number of learnable parameters for each best model configuration can be found in Table 7.
@@
Training and evaluation. Five splits are generated for each dataset, with 50%/25%/25% of the data going to the training, validation, and test sets, respectively; the exception is ZINC, for which the predefined splits are used (Irwin et al., 2012). The optimal hyperparameter configuration is chosen by selecting the best average performance over the five validation sets (details in Appendix C.2). One performance metric is reported per dataset. Specifically, predictive accuracy is used for Cora, Citeseer, PubMed, Amazon, Roman Empire, MUTAG, PROTEINS, NCI1, NCI109, IMDB-BIN, IMDB-MUL, and REDDIT; AUC-ROC is used for Minesweeper and Tolokers; mean squared error (MSE) is used for election, bachelor, birth, death, income, migration, and unemployment; and mean absolute error (MAE) is used for ZINC. For each dataset, the mean and standard deviation of the chosen metric are computed across the five test sets and reported in Table 1 (where OOM stands for ‘out of memory’).
@@ 5.2 Main Results
As seen from Table 1, higher-order neural networks (based on hypergraphs, simplicial, and cell complexes) achieve the best performance on fifteen of twenty-two datasets, whereas GNNs achieve the best performance on six datasets, and tie on the Unemployment dataset. GNNs perform best on node regression in the majority of cases (five out of seven). These best results obtained by GNNs are closely matched by TNNs, since the latter achieve metrics within one standard deviation from the former. In contrast, in nine out of sixteen datasets TNNs outperform GNNs, and attain performance metrics that are higher by more than one standard deviation with respect to GNNs. In other words, in situations where higher-order networks outperform GNNs, the performance gap is more pronounced. It is also noted that, for demonstration purposes, only one fixed lifting is considered to transform graph data to each of the considered topological domains (see Appendix C.4). These results suggest that TNNs have an advantage over GNNs, even without lifting optimization. More importantly, for the context of this paper, the benchmarks demonstrate the degree of comparisons that can be performed with TopoBench across models and datasets.
@@ 5.3 Ablation Study
This ablation study examines how different readout strategies influence performance in neural networks built on higher-order domains, highlighting the importance of node-level signal updates and pooling choices. First, graph and hypergraph (neural network) models differ from simplicial and cell complex (neural network) models in terms of the domains and, subsequently, representations they support. Graph and hypergraph models can output two types of representations: node representations and edge or hyperedge representations. In contrast, the output of simplicial and cell models depends on the different types of cells present (00-cell up to $n$-cells) and on the model itself. For example, a simplicial or cell complex model may process an $n$-cell input but may not produce an $n$-cell output. The backbone_wrapper in TopoBench addresses these differences in the underlying domains of the models.
@@
There is a second difference, which is inherent in the TNNs themselves. Consider a downstream classification task. For graphs, the standard practice is to perform classification over pooled node features. However, this aspect has not been extensively studied in the TDL literature. For instance, a simplicial or cell model may update 1-cell representations (edges) or 2-cell representations (cycles or triangles) while leaving 0-cell (node) representations unchanged, making direct pooling over nodes potentially ineffective. One could consider more elaborate update processes in which different $n$-cell representations are combined, but this renders pooling more intricate for higher-order domains. These architectural considerations are complex and remain open research questions in TDL.
@@
Nevertheless, to fairly compare different neural network architectures, this second difference must be addressed. To that end, this ablation study considers two types of readouts to enable a rigorous evaluation: direct readout (DR), where the downstream task is performed directly over the 00-cell representation, and signal down-propagation (SDP), where information from higher-order cell representations is iteratively fused down to 00-cell representations using appropriate incidence matrices, followed by a linear projection over the concatenated $(n-1)$-cell signal and the fused $(n-1)$-cell representation. For instance, if a simplicial or cell complex model outputs 00-cell, $1$-cell, and $2$-cell representations, the signal propagates from $2$-cells to $1$-cells and then from $1$-cells to 00-cells during readout. The downstream task is then performed over the updated 00-cell representations. [^4] See Appendix A.3 for an introduction to TNNs and Higher-Order Message Passing on topological domains.
@@
Table 2 shows that the best-performing readout type depends on how a model propagates signals internally. For example, the CWN model does not update 00-cell representations, so the SDP strategy performs notably better. Conversely, CCCN, SCCNN, and SCN propagate information to 00-cells, making SDP readout yield only small or negligible changes in performance. Further details are available in Appendix C.4.
@@
These results underscore the impact of structural properties in TNNs. The performance variations observed in this ablation study emphasize the critical role of architectural and lifting decisions for higher-order learning models. By enabling comparisons across a wide range of models and datasets, TopoBench facilitates deeper insights and drives advancements in TDL.
@@ 6 Concluding Remarks, Limitations, and Future Work
This paper has introduced TopoBench, an open-source benchmarking framework for TDL. By organizing the TDL pipeline into a sequence of modular steps, TopoBench simplifies the benchmarking process and accelerates research. A key feature of TopoBench is its ability to map graph topology and features to higher-order topological domains such as simplicial and cell complexes, enabling richer data representations and more detailed analyses. In addition, TopoBench provides direct access to a wide variety of real and synthetic datasets, covering both graph-based and higher-order domains. The effectiveness of TopoBench has been demonstrated by benchmarking several TDL architectures across diverse learning tasks and datasets, offering insights into the relative advantages of different models.
While TopoBench already addresses several challenges in TDL, it also has limitations that point to promising directions for future enhancements. One area is the integration of learnable liftings, which are currently not implemented. This direction could enable task-specific topological representations learned dynamically from data. A second limitation lies in the broader scarcity of standardized, real-world higher-order datasets. Although TopoBench incorporates numerous datasets for hypergraph, simplicial, and cell complexes, this remains an active area of expansion. Providing more built-in higher-order datasets will further streamline research in TDL. Finally, extending the set of evaluation metrics beyond classification or regression accuracy to include more TDL-specific measures of expressivity, explainability, and fairness (Papamarkou et al., 2024) is another avenue of growth.
Moving forward, the modular design of TopoBench invites contributions from the community. Researchers and practitioners are encouraged to contribute to TopoBench by introducing new learnable liftings, adding datasets, and developing specialized performance metrics. These efforts will not only strengthen the benchmarking ecosystem of TDL but also help drive innovation in topological deep learning more broadly.
Code Availability and Reproducibility
@@
The code for TopoBench is publicly available on GitHub under the MIT license: https://github.com/geometric-intelligence/TopoBench. The codebase employs continuous integration, is fully documented, and provides comprehensive API documentation at https://geometric-intelligence.github.io/topobench/index.html.